With a tail wind*, a helicopter traveled 300 mi in 1 hr 40 min. The return trip against the same wind took 20 min longer. Find the wind speed* and also the air speed* of the helicopter.

*air speed- the speed of an aircraft in still air
*wind speed- the speed of the wind
*tail wind- a wind blowing in the same direction as the path of the aircraft.

Answer by checkley71(8403)

300/100=3 OR 3 MILES PER MINUTE=180 MPH
300(120)=2.5 0R 2.5 MILES PER MINUTE OR 150 MPH.
180-W=150+W
-W-W=150-180
-2W=-30
W=-30/-2
W=15 MPH FOR THE WIND.
PROOF
300/180=1 2/3 HOURS = 1 HOUR & 40 MINUTES.
300/150=2 HOURS
2 HRS-(1 HOUR & 40 MINUTES)=20 MINUTES LONGER FLIGHT.